b1az3:

The data list shows the cost, in dollars, of different brands of sneakers at a shoe store. 44.99, 59.99, 32.30, 65.00, 52.59 What is the population standard deviation, to the nearest tenth, of the data (assuming this is all of the data)?

AZ:

You're really asking the same question.... a week later...? Even after I respond multiply times... But you won't even respond???

AZ:

\(\sigma =\sqrt{ \dfrac{\sum(x_i-\mu)^2}{N}}\) where \(\sigma\) = population standard deviation \(\sum\) = the sum from 1 to N \( x_i\) = each individual data point \(\mu\) = mean of the data list N = total number of data points so to first find the population standard deviation, can you tell me what the average or mean is? 44.99, 59.99, 32.30, 65.00, 52.59 How many numbers do we have? That's going to be our N. Add them all and divide by that number N. That will be our mean. What do you get?

AZ:

Just answer my questions and we can find the standard deviation one step at a time How many numbers are in your list: 44.99 59.99 32.30 65.00 52.59 Can you count how many numbers there are?

AZ:

Then add all those numbers up 44.99 + 59.99 + 32.30 + 65.00 + 52.59 What do you get? And then divide the number that you get now by how many numbers there were in your list

AZ:

That's going to give us your \(\mu\) or mean Once you answer those few questions I asked you above, we can move to the other variables and complete your question :)

b1az3:

1 attachment
AZ:

But like if you're seriously that incompetent and really don't want to work through the question with me... have you even decided to put any effort into your question? Or like even search up a calculator https://www.easycalculation.com/statistics/standard-deviation.php

b1az3:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Just answer my questions and we can find the standard deviation one step at a time How many numbers are in your list: 44.99 59.99 32.30 65.00 52.59 Can you count how many numbers there are? \(\color{#0cbb34}{\text{End of Quote}}\) 6

AZ:

Do you want to count them again for me? Maybe try using your hands this time??

b1az3:

5

AZ:

Much better

AZ:

Now add all the numbers 44.99 + 59.99 + 32.30 + 65.00 + 52.59 =

b1az3:

254.87

snowflake0531:

I suggest dividing that by 5.

AZ:

Divide by 5

b1az3:

50.974

b1az3:

= 51

AZ:

keep it as 50.974

b1az3:

it says to the nearest tenth,

AZ:

so now we do that \( x_i - \mu\) part so what is 44.99 - 50.974 = 59.99 - 50.974 = 32.30 - 50.974 = 65.00 - 50.974 = 52.59 - 50.974 =

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @b1az3 it says to the nearest tenth, \(\color{#0cbb34}{\text{End of Quote}}\) #1 51 isn't even the nearest tenth. That's a whole number #2 You only round once you get to the last final answer. Not any time sooner than that

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ so now we do that \( x_i - \mu\) part so what is 44.99 - 50.974 = 59.99 - 50.974 = 32.30 - 50.974 = 65.00 - 50.974 = 52.59 - 50.974 = \(\color{#0cbb34}{\text{End of Quote}}\) so tell me what you get for each one of these because we still have quite a ways to go to get to your final answer